OpenStudy (anonymous):

Martin draws a card from a standard deck of cards, does not replace the card, and draws another card. What is the probability he draws an ace card and a queen?

Directrix (directrix):

Probability is the ratio of the number of favorable outcomes to the number of possible outcomes. There are 52 cards in the deck. 4 are aces and 4 are queens. Draw one card. P(Ace) = 4/52 The first card is not replaced. That leaves 51 cards. P(Queen) = 4/51 P(Ace and Queen) = 4/52 * 4/ 51 = ? @Raylynngrace